OpenStudy (anonymous):

Input the equation for the following line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. A line with y-intercept of 2 that is parallel to 2x + y = -5.

OpenStudy (anonymous):

Well, you know that for a line to be parallel it has to have the same slope. If you solve the equation for y you see that: y=-2x-5 So the slope is -2. Using slope-intercept form (y=mx+b) you can write your new line as: y=-2x+2 (the -2 we got from the fact its parallel, and the 2 is the y-intercept) Then just rearrange it to standard form. y+2x=2

OpenStudy (anonymous):

could you help me with one more?

OpenStudy (anonymous):

Sure :P

OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (-8, 0), (1, 5)

OpenStudy (anonymous):

Okay, firstly, you know that the slope of a line is: \[m=\frac{\Delta y}{\Delta x}\] So your change in y would be: 5-0=5 Your change in x would be 1-(-8)=9 So your slope is: \[\frac{\Delta y}{\Delta x}=\frac{5}{9}\] Then you need to use point slope form: \[y-y_0=m(x-x_0)\] Either point will work. \[y-5=\frac{5}{9}(x-1) \rightarrow y-5=\frac{5x}{9}-\frac{5}{9}\] Adding 5 to the right gives you: \[y=\frac{5x}{9}+\frac{40}{9}\] Then you want to put the x on the left side giving: \[\frac{-5x}{9}+y=\frac{40}{9}\] And if you wanted to be really clever you could multiply both sides by 9 to get ride of the denominator giving: \[-5x+9y=40\] Just to make it look nice. :P

OpenStudy (anonymous):

could you help me with one more ? :)

OpenStudy (anonymous):

Yeah go for it. I'm jumping around a few questions so it might take me a moment.

OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (anonymous):

dont forget about me!

OpenStudy (anonymous):

For this one you want to do the same thing. You want to find the change in y and x. The change in y is -8-(-3)=-5 And the change in x is 4-7=-3 So your slope is 5/3. Using point slope again you have: y-(-3)=(5/3)(x-7) \[y+3=\frac{5x}{3}-\frac{35}{3}\] Subtracting 3 gives you: \[y=\frac{5x}{3}-\frac{44}{3}\] Putting the x on the left side and multiplying by 3 to get: \[3y-5x=-44\]

OpenStudy (brooke2353):

y = ax + b -3 = 7a + b b = -7a - 3 and -8 = 4a + b b = -4a - 8 then -7a - 3 = -4a - 8 3a = 5 a = 5/3 b = - 4(5/3) - 8 b = - 20/3 - 24/3 = -44/3 y = (5/3)x - 44/3 3y = 5x - 44 5x - 3y = 44